OpenStudy (anonymous):

-6f+13=2f-11 please work the problem out! Thank you sooo much!!

OpenStudy (anonymous):

what happens when you add 11 to both sides?

OpenStudy (anonymous):

You don't. You subtract 2 from both sides .

OpenStudy (anonymous):

ok, what happens when you subtract 2 from both sides?

OpenStudy (anonymous):

You get -8+13=-11

OpenStudy (anonymous):

ok, but you forgot to write the F in there somewhere

OpenStudy (anonymous):

I meant to put the f behind the 8. But It has an f on my paper.

OpenStudy (anonymous):

ok, you got all the Fs together, now get the other numbers together move that 13 to the other side

OpenStudy (anonymous):

Draw it

OpenStudy (anonymous):

you have so far: -8f + 13 = 11 so to get the numbers together, subtract 13 from both sides

OpenStudy (anonymous):

So you would add 11 and 13, right?

OpenStudy (anonymous):

-8f + 13 - 13 = 11 - 13

OpenStudy (anonymous):

what is 13-13 and what is 11-13 ?

OpenStudy (anonymous):

I got f=3

OpenStudy (anonymous):

good work... easy as that

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

2z-31=-9z+24 ????